OpenStudy (anonymous):

A right square pyramid has base perimeter 62.4m and height 6.4m. Calculate the volume of the pyramid to the nearest cubic metre.

OpenStudy (anonymous):

1/3 * area of base* height=1/3 *(62.4/4)^2* 6.4 cubic metre= 519.168 cubic metre